OpenStudy (jazzyluvsu123):

Checking my work plz tell me if it is right 1.5.7+0.98=6.68 Fraction--> 6/21+7/21= 13/21 (Can it be reduced) 4. 54.86 (divided by) 12=4.57 1/6 <-- fraction @mathstudent55 @agent0smith

OpenStudy (agent0smith):

5.7+0.98=6.68 yes but you could've used a calculator 6/21+7/21= 13/21 13 is a prime number 54.86 (divided by) 12=4.57 1/6 why is there a fraction AND a decimal...?

OpenStudy (jazzyluvsu123):

I don't I had a remander and couldn't remember what to do with it ? what should I do @agent0smith

OpenStudy (agent0smith):

You just keep dividing... idk what you're doing with it... long division?

OpenStudy (jazzyluvsu123):

I got it thnks